EVA Report – February 5th

Crew 221 EVA Report 05 Feb 2020

EVA # 1

Author: Maria Grulich (Commander)

Purpose of EVA: Collection of regolith and geological samples to be used as an ISRU growth substrate for some botanical experiments in the GreenHab, and geology and microbiology experiments.

Start time: 09:35 am

End time: 12:30 pm

Narrative: Today we planned our first long EVA. Before going on EVA, crew commander M.Grulich took a sample of the bacteria on the space suits, in particular, the cleaned EVA suit and the suit after coming back to the Hab to see how contaminated the space suits are getting during an EVA. L.Monge, J.Todd and M.Grulich went out to drive towards the Moons with the aim to collect samples for ISRU and microbiology projects. The Martian weather is currently very much in our favor. The sun was smiling at us and there was almost no wind. M.Grulich took the lead on spirit to guide the team towards the Moons. Once near the Moons, the EVA crew left the rovers along Cow Dung Road. Unfortunately, the EVA preparations took one hour longer than expected therefore the first planned stop at Cowboy Corner was canceled. The EVA team did not manage to identify Brahe Highway 1572 therefore the team probably stopped near Factory Butte View and turned left into the field.

Once at the Moons, crew engineer Monge identified the first out of four locations to collect soil for the ISRU project to grow black corn in it. Crew Commander M.Grulich was looking for geology traces where in former times water could have run through to find and cultivate microbes. One location seems to be very promising – the soil showed bubbles which indicates bacteria activities. While taking samples for microbiology, the team was very careful using extra gloves on top of the EVA gloves and sterile sample containers and little shovels that have had been cleaned prior to the EVA. Also, one air sample was taken to see if there are any microbes in the Martian air around the Moon area. J. Todd was also taking geology samples to investigate any traces of salt indicating former water on the planet. All objectives of the EVA have been accomplished. Shortly before returning to the airlock the EVA team, was checking on the weather station installed yesterday and observed that the water was frozen, making it impossible to read out measurements. This was reported to the support team in the habitat to take action in the second EVA of that day.

The EVA team was working efficiently and very well together. A secondary objective of each EVA is to trace the actual taken road and average heart rate. Unfortunately after 20min, the tracking device stopped recording, which will be investigated in further EVAs to analyze and improve future EVA preparation.

Destination: Gray Moon, Beige Moon, The Moon

Coordinates (use UTM NAD27 CONUS): 4254500/516500 (Gray Moon), 4254350/516500 (Beige Moon), 4254800/515700 (The Moon)

Participants: Maria Grulich (Commander), Luis Monge (Crew Engineer), Jessica Todd (GreenHab Officer)

Road(s) and routes per MDRS Map: Cow Dung Road (0110) parked near Factory Butte View walked into the field.

Mode of travel: Driving and Walking

Vehicles used: Curiosity and Spirit

Copyright © The Mars Society. All rights reserved. | Main Site